You decide to take advantage of the current online dating craze and start your own Web site. You know that you have 450 people who will sign up immediately and, through a careful marketing research and analysis, determine that membership can grow by 25 percent in the first two years, 21 percent in year 3, and 18 percent in Year 4. How many members do you expect to have at the end of four years?
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

A=450*(1.25)^2*(1.21)*(1.18)

=450*2.2309375

=1004(Approx).